Joey’s Bike Shop sells new and used bicycle parts. Although a majority of its sales are cash sales, it makes a significant amount of credit sales. During 2018, its first year of operations, Joey’s Bike Shop experienced the following:

Sales on account $ 250,100
Cash sales 646,200
Collections of accounts receivable 237,595
Uncollectible accounts charged off during the year 1,199

Required

Assume that Joey's Bike Shop uses the allowance method of accounting for uncollectible accounts and estimates that 1 percent of its sales on account will not be collected. Answer the following questions:

  1. What is the Accounts Receivable balance at December 31, 2018?
  2. What is the ending balance of Allowance for Doubtful Accounts at December 31, 2018, after all entries and adjusting entries are posted?
  3. What is the amount of uncollectible accounts expense for 2018?
  4. What is the net realizable value of accounts receivable at December 31, 2018?
Answer #1
a
Accounts Receivable balance at December 31, 2018 11306 =250100-237595-1199
b
Allowance for Doubtful Accounts at December 31, 2018 1302 =(250100*1%)-1199
c
Uncollectible accounts expense for 2018 2501 =(250100*1%)
d
Net realizable value of accounts receivable 10004 =11306-1302